OpenStudy (anonymous):

how to do absolute value equations 1. 5I3x-4I=x+1 2.3Ix+6I=9x-6

OpenStudy (anonymous):

FIRST ONE: \[1.5\left| 3x-4 \right|=x+1\] Split it up: 1.5(3x-4)=x+1 OR 1.5(4-3x) = x+1 4.5x-6 = x+1 OR 6-4.5x =x+1 3.5x = 7 OR 5=5.5x x={2,1.1}

OpenStudy (amoodarya):

if its 3Ix+6I=9x-6 ?! has one root in x=4 1st simplify by 3 then do like "enelius" said

OpenStudy (amoodarya):

by solving 2nd as I say you get x=4, x=-1 x=4 is correct but x=-1 is not (check in equation )